# Exploiting Hatred by Targets for Hate Speech Detection on Vietnamese Social Media Texts

### v1

@@ 1 Introduction
There are several attempts at the hate speech problem for Vietnamese social networks, such as the VLSP-HSD Shared Task 2019 [3], the ViHSD [4] and the ViCTSD [5] datasets. These works are valuable in constructing and evaluating the computer machine learning model to automatically identify toxic comments on social media space. Besides, the author in [6] introduces a solution for hate speech detection on social networks by the streaming approach using the online streaming platform, which helps the system process for real-time processing. However, the research just focuses on the classification task, which means the output that is hate or not hate is based on the whole comments. According to [7], it is difficult to capture the abuse and hateful content because users tend to replace the harmful words with more benign phrases and use subtle forms to prevent the detection. Thus, the challenge for the hate speech detection model is to exploit the hateful content, including not only the hateful action represented in the text but also the aimed target to increase the accuracy of classification models. In addition, measuring the public attitudes toward a specific target through the comments from social network media will help to understand more about the social phenomenon, and help solve the issues in social science [8]. Moreover, from the output people not only to find whether these comments are hate speech or not but also want to interpret the specific targets that hateful comments are aimed at. Therefore, in this paper, we provide new datasets named ViTHSD that help to interpret hate speech in the comments.
@@ 2 Related Works
Dataset Labels Total Size Target Labels Rationales
VLSP-HSD (2019) [3] Hate Speech, Offensive, Normal 20,345 No No
ViHSD (2021) [4] Hate Speech, Offensive, Normal 33,400 No No
ViHOS (2023) [13] Hate Speech, Offensive, Normal 11,065 No Yes
ViTHSD (Ours) Labels for five different aspects 10,000 Yes No
@@ 3.1 Annotation Procedure
Annotation guidelines: The ViTHSD dataset comprises four labels, namely HATE, OFFENSIVE, CLEAN, and NORMAL, pertaining to five different targets. Even though a comment may contain multiple targets in ViTHSD, the degree of hate or abuse involved can vary for each target. Table 2 describes the definitions and examples of each hatred level. The principal objective of the annotation process is to identify the groups that are being targeted by hate speech in the comments and determine the extent of hostility directed towards those communities. To ensure that the labeling process is facilitated, we decided to base the selection of target groups on the characteristics of the dataset. We included five targets based on individuals (1), groups (2), religion/creed (3), race/ethnicity (4), and politics (5). The details of those targets are described below:
  • Individualss which is Hate speech based on the identities of individuals can be attacked on elements of physical disability, sexual orientation, etc. This is one of the most popular attack methods on Vietnamese social networks, often expressed in sentences containing names or personal pronouns in the Vietnamese language. According to a report by Vietnam Digital 2021, 53% of Vietnamese internet users have experienced online harassment, with 21% of them experiencing hate speech [4]. These types of personal attacks are commonly referred to as cyberbullying and cyberaggression [25]. For example:
@@
NORMAL Comments with content that is not related to the targets under consideration. There is no need to consider the comments containing offensive words, hate or not. Example 1: Nhanh thực sự (English: It is really fast). This comment is thoroughly clean, contains no offensive or hate words, and attacks no one.
@@
Example 3: Đi tìm nó đánh chết mẹ nó luôn chị (English: Go find him and kill his mother too, sister.). The phrase kill his mother appears in this comment. This is an inflammatory phrase that can cause hatred and has the potential to threaten to cause anyone to feel unsafe. Example 4: Bán Nam bán Nữ (English: semi-male, semi-female.). No words expressing sarcasm or hatred, but content implies insulting targets, specifically people’s sexual characteristics.
@@ 3.2 Dataset Creation Steps
The dataset creation process began with a pilot annotation round, followed by the main annotation task. The main annotation task was inspired by the work of Mathew et al. [20]. Inter-annotator agreement (IAA) was computed using the Cohen Kappa index ($k$) [29], and the final labels were selected using a majority voting algorithm [30]. Figure 5 displays the inter-annotator agreement between each annotating pair for each target. The dataset was generated using a two-stage process, as depicted in Figure 1.
@@ 3.3 Inter-annotator agreement
Figure 2: The confusion matrix between the annotators in a set of 5000 comments in each target without levels, computed by the Cohen kappa index ($k$)
@@
Table 3: Cohen’s Kappa ($k$) average for each aspect, without levels.
@@
Table 4: Average Cohen’s Kappa ($\overline{k}$) for each target have levels
@@
We found significant variation in the average Cohen’s Kappa results for each aspect with levels. Table 4 shows the inter-annotator agreement analysis results for a dataset of 5,000 comments, indicating moderate agreement among annotators. The Race/Ethnicity target had the highest level of agreement, while Religion/Creed had the lowest. We also analyzed a larger dataset of 10,000 comments, which showed higher agreement levels among annotators than the smaller dataset. However, for the larger dataset, individuals and Religion/Creed targets had lower agreement scores compared to Group and Race/Ethnicity. Figures 4 and 5 provide clearer consensus between annotator pairs.
Figure 4: The Cohen’s kappa index ($k$) was used to calculate the confusion matrix between the annotators on 5000 comments per target.
Figure 5: The Cohen’s kappa index ($k$) was used to calculate the confusion matrix between the annotators on 10000 comments per target.
Analyzing the inter-annotator agreement shows that the annotators agree to a high degree when determining the aspect of a sentence. This suggests a shared understanding of the underlying themes commented upon. However, there was significant variability between annotators when assessing the level of hate within a given aspect. This variability was particularly evident for the Religion/Creed and Race/Ethnicity targets, suggesting that the perception of hate speech related to these two targets is more subjective and less consensual among annotators. These findings highlight the complexity of identifying and assessing hate speech, especially when the targets are related to sensitive issues such as ethnicity and religion.
@@ 3.4 Dataset Analysis
Table 5 illustrates the general statistics about the ViTHSD dataset. The vocabulary size is calculated by token level, and the average length is calculated by the total length of comments divided by the number of comments in each set. From Table 5, it can be seen that the train set has the biggest vocabulary size. However, the development (dev) set has the highest average length value, which means that the comments in the development set seem longer than the train and dev sets. In contrast with the average length, the vocabulary size in the dev set is lower than the two remains.
@@ 4.1 Baseline Models
We follow the system architecture in [22] to construct the baseline model for the targeted hate speech detection task. Figure 7 illustrates the fine-tuned BERT model for this task. According to the original work [22], we unfreeze the four last layers of BERT for training parameters. Then, the parameters from the BERT are passed to the dropout layer and spread the probability to the five Dense layers. These five Dense layers correspond to five targets in the dataset, including Individual, Group, Religion, Race, and Politics. On each target, there are four levels, including Normal, Clean, Offensive, and Hate (as described in Section 3). The predicted targets and relevant levels are which have the highest probability. To adapt the BERT for the Vietnamese language model, we used the pre-trained multilingual and monolingual BERTology as shown in Table 11.
@@
In addition, to enhance the ability of the model to extract valuable features from the texts for hate speech detection, we combine the Bi-GRU-LSTM-CNN [9] with the pre-trained BERTology. Figure 8 shows our proposed model architecture for targeted hate speech detection, which is a modification of the original Bi-GRU-LSTM-CNN model [9]. First, we replace the embedding layer with the pre-trained BERT layers to take advantage of BERT as a robust language model for text representation. Then, for the last layer, we use five Dense layers instead to treat the distribution of probability within five targets. For each target, there are four levels of hate speech.
For the pre-trained BERTology models, as mentioned in Table 11, we choose the model which has the highest performance to construct the prediction model for the online streaming system (as described in Section 4.2). Our empirical results of each model are illustrated in Section 5.
@@ 4.2 Online Streamming System
Various real-time data processing methods and systems have been introduced, such as the data streaming process by Nagarajan et al.[40] and the real-time processing of Twitter tweets using Spark Streaming [41]. Rengarajan et al. [42] has also presented a streaming pipeline design for big data with Kafka [43]. Studies have been done on real-time streaming systems to detect hate speech in Vietnamese. For example, Long-An Doan et al. [24] used machine learning and big data technology to detect hate speech in Vietnamese YouTube comments. Another study, by Khanh Q. Tran et al. [23], developed a novel hate speech detection model that combines a pre-trained PhoBERT model and a Text-CNN model and uses EDA techniques to improve classification performance. To sum up the paradigm for constructing a system that can process on real-time stream of data, Dinh et al. [44] proposed a scalable framework that includes four main components. First is the crawling modules that collect real-time text comments from social networks. Then, these text comments are pre-processed to create clean text. Next, the cleaned data is fitted to the classification model, which has been trained and evaluated before, to make the prediction labels. Finally, to deliver valuable information to the end-users, the prediction results are visualized on a real-time dashboard.
@@ 6.1 Execution Evaluation on Streaming Data
We investigate the performance of XLM-R-Bi-GRU-LSTM-CNN and the ViSoBERT-R-Bi-GRU-LSTM-CNN on a streaming system to measure the efficiency of both models in a practical real-time application. We integrate two models into the streaming systems to predict about 1,000 comments coming from YouTube live streaming in real time. Then we measure the time execution for predicting each comment of each model. Table 14 shows the statistical information in time execution for each model, and Figure 11 illustrates the distribution of time execution for each comment of both XLM-R-Bi-GRU-LSTM-CNN and ViSoBERT-R-Bi-GRU-LSTM-CNN models. We run the experiments on the server with 16GB RAM, 50GB Disk SSD SATA 3, and 4 CPU Intel Core i5 (i5-8250U).
Table 14: Statistical information about prediction times and size of classification models on the streaming system. The time unit is milliseconds, and the size is Gigabyte (Gb)
@@ 7 Conclusion and Future Work
In this paper, we have built a new dataset called ViTHSD to detect targeted hate speech in Vietnamese texts. The dataset comprises 10,000 comments manually annotated by humans across five targets: Individual (1), Group (2), Religion/creed (3), Race/Ethnicity (4), and Politics (5). Each target includes four levels of hate: NORMAL, CLEAN, OFFENSIVE, and HATE, except NORMAL which indicates no specific target. We implemented a comprehensive annotation process, from annotator selection to consensus evaluation methodology. The ViTHSD dataset has a moderate level of inter-annotator agreement, measured by the Cohen-Kappa K-average metric ($k$=0.45). We assess multiple aspects of hate speech detection by testing several state-of-the-art models on this dataset. The XLM-R Bi-GRU-LSTM-CNN model achieves the best performance, with a Macro F1 score of 72.01% for the Target task only and 51.74% for the Target task with a level. Besides, the ViSoBERT - a monolingual pre-trained model on Vietnamese social texts, obtained better performance than XLM-R with fewer parameters and less time execution. Additionally, we have successfully constructed a streaming architecture that runs smoothly and demonstrated its functionality through a continuous 60-minute video stream.
Challenges such as vocabulary inconsistencies in social media texts, which include diverse languages, acronyms, slang, and mixed codes, result in inaccurate predictions. Addressing low-accuracy models will be a significant challenge going forward. Our error analysis revealed that developing a system capable of recognizing offensive language on social networks in Vietnam is incredibly challenging. Furthermore, a shortage of resources to evaluate the system at a higher level presents a significant obstacle for us. We will persist in refining the dataset’s quality while investigating system upgrades that can meet real-world application demands.
@@
The dataset will be published for research purposes. Please contact the corresponding author for data inquiry.

### v3

# ViTHSD: Exploiting Hatred by Targets for Hate Speech Detection on Vietnamese Social Media Texts

@@ 1 Introduction
There are several attempts at the hate speech problem for Vietnamese social networks, such as the VLSP-HSD Shared Task 2019 [3], the ViHSD [4] and the ViCTSD [5] datasets. These works are valuable in constructing and evaluating the computer machine learning model to automatically identify toxic comments on social media space. Besides, the author in [6] introduces a solution for hate speech detection on social networks by the streaming approach using the online streaming platform, which helps the system process for real-time processing. However, the research just focuses on the classification task, which means the output that is hate or not hate is based on the whole comments. According to [7], it is difficult to capture the abuse and hateful content because users tend to replace the harmful words with more benign phrases and use subtle forms to prevent the detection. Thus, the challenge for the hate speech detection model is to exploit the hateful content, including not only the hateful action represented in the text but also the aimed target to increase the accuracy of classification models. In addition, measuring the public attitudes toward a specific target through the comments from social network media will help to understand more about the social phenomenon, and help solve the issues in social science [8, 9]. Moreover, from the output people not only want to find whether these comments are hate speech or not but also want to understand the specific targets that hateful comments are aimed at. According to [10], this is called interpretability, which means ”humans can understand the cause of a decision”. The interpretability of the specific target in the hate speech can help mitigate the bias features in the users’ text for hate speech detection [11]. Therefore, in this paper, we provide new datasets named ViTHSD that help to interpret hate speech in the comments.
@@ 2 Related Works
Dataset Labels Size Task Target Labels Rationales
VLSP-HSD (2019) [3] Hate, Offensive, Normal 20,345 Multiclass classification No No
ViHSD (2021) [4] Hate, Offensive, Normal 33,400 Multiclass classification No No
ViCTSD [5] Toxic and Non-toxic 10,000 Binary classification No No
ViHOS (2023) [16] Spans determining the hate speech pieces in the comments 11,065 Span labeling No Yes
ViTHSD (Ours) Hate, Offensive, Normal corresponding to each of five targets 10,000 Multilabel classification Yes No
@@ 3.1 Annotation Procedure
Targets in Hate speech detection: According to [2], although the hate speech definition mentions various aspects to detect the hate speech, the general hate speech on social media focuses on some specific targets such as Sexual orientation, Race, and Religion. Other targets such as Gender and Disability are very few thus they have less meaning when separated. Besides, the authors in [28] show various types of hate according to different aspects such as culture, economics, crimes, rapism, women oppression, history, and other/generic types. However, some hate speech appears in complex linguistic styles (e.g. indirect sarcasm and humor), making it hard to mitigate the bias among targets. Referred from the previous works in [2, 23, 29, 30], we proposed a list of targets that mostly appear in Vietnamese comments including Groups, Individuals, Religion, Race, and Politics. These targets play the role of exploiting the subject expressed in the comments by users, which helps efficiently detect hate speech by targets. The details about these targets are described in the Annotation guidelines section.
Annotation guidelines: The ViTHSD dataset comprises four hatred labels, namely HATE, OFFENSIVE, CLEAN, and NORMAL, pertaining to five different targets. Even though a comment may contain multiple targets in ViTHSD, the degree of hate or abuse involved can vary for each target. Table 2 describes the definitions and examples of each hatred level. There are three hatred levels, which are CLEAN, OFFENSIVE, and HATE. The NORMAL is not the hatred level since it indicates that the target is not mentioned. The principal objective of the annotation process is to identify the groups that are being targeted by hate speech in the comments and determine the extent of hostility directed towards those communities. To ensure that the labeling process is facilitated, we decided to base the selection of target groups on the characteristics of the dataset. We included five targets based on individuals (1), groups (2), religion/creed (3), race/ethnicity (4), and politics (5). The details of those targets are described below:
  • Individuals which is Hate speech based on the identities of individuals can be attacked on elements of physical disability, sexual orientation, etc. This is one of the most popular attack methods on Vietnamese social networks, often expressed in sentences containing names or personal pronouns in the Vietnamese language. According to a report by Vietnam Digital 2021, 53% of Vietnamese internet users have experienced online harassment, with 21% of them experiencing hate speech [4]. These types of personal attacks are commonly referred to as cyberbullying and cyberaggression [31]. For example:
@@
NORMAL (*) Comments with content that is not related to the targets under consideration. There is no need to consider the comments containing offensive words, hate or not. Example 1: Nhanh thực sự (English: It is really fast). This comment is thoroughly clean, contains no offensive or hate words, and attacks no one.
@@
Example 4: Đi tìm nó đánh chết mẹ nó luôn chị (English: Go find him and kill his mother too, sister.). The phrase kill his mother appears in this comment. This is an inflammatory phrase that can cause hatred and has the potential to threaten to cause anyone to feel unsafe.
Example 5: Bán Nam bán Nữ (English: semi-male, semi-female.). No words expressing sarcasm or hatred, but content implies insulting targets, specifically people’s sexual characteristics.
(*) NORMAL is not the hatred level because it expresses that the target is not mentioned. There are three hatred levels including CLEAN, OFFENSIVE, and HATE.
@@
Figure 1: An example of annotated data by users on Google sheet.
Annotation tool: We used the Google sheet [^1] https://docs.google.com/spreadsheets for annotating the data. Each annotator annotates their label on their own sheet as shown in Figure 1. Each comment (one comment per row) has five columns corresponding to five targets. The annotator assigns one numeric value denoting the hatred level for each target as 1 for clean, 2 for offensive, and 3 for hate (see Table 2 for details about the level of hatred level). If the comment does not mention the target, the annotator assigns the value 0 as the normal label (as shown in Table 2).
@@ 3.2 Dataset Creation Steps
The dataset creation process began with a pilot annotation round, followed by the main annotation task. The main annotation task was inspired by the work of Mathew et al. [23] and Luu et al. [4]. Inter-annotator agreement (IAA) was computed using the Cohen Kappa index ($k$) [35], and the final labels were selected using a majority voting algorithm [36]. Figure 6 displays the inter-annotator agreement between each annotating pair for each target. The dataset was generated using a two-stage process, as depicted in Figure 2.
@@ 3.3 Inter-annotator agreement
Figure 3: The confusion matrix between the annotators in a set of 5,000 comments in each target without levels, computed by the Cohen kappa index ($k$)
@@
Table 3: Cohen’s Kappa ($k$) average for each aspect without levels.
@@
Table 4: Average Cohen’s Kappa ($\overline{k}$) for each target with levels
@@
We found significant variation in the average Cohen’s Kappa results for each aspect with levels. Table 4 shows the inter-annotator agreement analysis results for a dataset of 5,000 comments, indicating moderate agreement among annotators. The Race/Ethnicity target had the highest level of agreement, while Religion/Creed had the lowest. We also analyzed a larger dataset of 10,000 comments, which showed higher agreement levels among annotators than the smaller dataset. However, for the larger dataset, individuals and Religion/Creed targets had lower agreement scores compared to Group and Race/Ethnicity. Figures 5 and 6 provide clearer consensus between annotator pairs.
Figure 5: The Cohen’s kappa index ($k$) was used to calculate the confusion matrix between the annotators on 5,000 comments per target.
Figure 6: The Cohen’s kappa index ($k$) was used to calculate the confusion matrix between the annotators on 10,000 comments per target.
Overall, analysis of the inter-annotator agreement shows that the annotators agree to a high degree when determining the aspect of a sentence. This suggests a shared understanding of the underlying themes commented upon. However, there was significant variability between annotators when assessing the level of hate within a given aspect. This variability was particularly evident for the Religion/Creed and Race/Ethnicity targets, suggesting that the perception of hate speech related to these two targets is more subjective and less consensual among annotators. These findings highlight the complexity of identifying and assessing hate speech, especially when the targets are related to sensitive issues such as ethnicity and religion.
@@ 3.4 Dataset Analysis
We divided our dataset into the training, development, and test sets with a proportion of approximately 70/10/20. Table 5 illustrates the general statistics about the ViTHSD dataset. The vocabulary size is calculated by token level, and the average length is calculated by the total length of comments divided by the number of comments in each set. From Table 5, it can be seen that the train set has the biggest vocabulary size. However, the development (dev) set has the highest average length value, which means that the comments in the development set seem longer than the train and dev sets. In contrast with the average length, the vocabulary size in the dev set is lower than the two remains.
@@ 4.1 Baseline Models
We follow the system architecture in [25] to construct the baseline model for the targeted hate speech detection task. Figure 8 illustrates the fine-tuned BERT model for this task. According to the original work [25], we unfreeze the four last layers of BERT for training parameters. Then, the parameters from the BERT are passed to the dropout layer and spread the probability to the five Dense layers. These five Dense layers correspond to five targets in the dataset, including Individual, Group, Religion, Race, and Politics. On each target, there are four levels, including Normal, Clean, Offensive, and Hate (as described in Section 3). We employ the Softmax function to compute the distribution probability to the four levels: normal, clean, offensive, and hate for each Dense layer. The predicted hatred levels corresponding to each target are which have the highest probability from each of the five Dense layers. If the target is predicted as ”Normal” level, it means the target is not mentioned in the comment. Otherwise, the remaining levels (Clean, Offensive, and Hate) determine the hatred intensity for each target. To adapt the BERT for the Vietnamese language model, we used the pre-trained multilingual and monolingual BERTology as shown in Table 11.
@@
In addition, to enhance the ability of the model to extract valuable features from the texts for hate speech detection, we combine the Bi-GRU-LSTM-CNN [12] with the pre-trained BERTology. The Bi-GRU-LSTM-CNN [12] employs 2 robust sequence processing models including Bi-GRU and Bi-LSTM to extract valuable features from the input sequence. Then it uses a CNN layer to fine-grained the information from Bi-GRU and Bi-LSTM before forwarding it to the Dense layer for classification. The Bi-GRU-LSTM-CNN model achieved optimistic results from the VLSP 2019 - Hate speech detection shared task [3]. Figure 9 shows our proposed model architecture for targeted hate speech detection, which is a modification of the original Bi-GRU-LSTM-CNN model [12]. First, we replace the embedding layer with the pre-trained BERT layers to take advantage of BERT as a robust language model for text representation. Then, for the last layer, we use five Dense layers instead to treat the probability distribution of hatred levels within five targets.
@@
Finally, for the pre-trained BERTology models as mentioned in Table 11, we choose the model which has the highest performance to construct the prediction model for the online streaming system (as described in Section 4.2). Our empirical results of each model are illustrated in Section 5.
@@ 4.2 Online Streamming System
Various real-time data processing methods and systems have been introduced, such as the data streaming process by Nagarajan et al.[46] and the real-time processing of Twitter tweets using Spark Streaming [47]. Rengarajan et al. [48] has also presented a streaming pipeline design for big data with Kafka [49]. Studies have been done on real-time streaming systems to detect hate speech in Vietnamese. For example, Long-An Doan et al. [27] used machine learning and big data technology to detect hate speech in Vietnamese YouTube comments. Another study, by Khanh Q. Tran et al. [26], developed a novel hate speech detection model that combines a pre-trained PhoBERT model and a Text-CNN model and uses EDA techniques to improve classification performance. To sum up the paradigm for constructing a system that can process a real-time stream of data, Dinh et al. [50] proposed a scalable framework that includes four main components. First is the crawling modules that collect real-time text comments from social networks. Then, these text comments are pre-processed to create clean text. Next, the cleaned data is fitted to the classification model, which has been trained and evaluated before, to make the prediction labels. Finally, to deliver valuable information to the end-users, the prediction results are visualized on a real-time dashboard.
@@ 6.1 Execution Evaluation on Streaming Data
We investigate the performance of XLM-R-Bi-GRU-LSTM-CNN and the ViSoBERT-R-Bi-GRU-LSTM-CNN on a streaming system to measure the efficiency of both models in a practical real-time application. We integrate two models into the streaming systems to predict about 1,000 comments from YouTube live streaming in real time. Then we measure the time execution (milliseconds) for predicting each comment of each model and summarize them in quartile information. Table 14 shows the statistical of time execution for each model, and Figure 12 illustrates the distribution of time execution for each comment of both XLM-R-Bi-GRU-LSTM-CNN and ViSoBERT-R-Bi-GRU-LSTM-CNN models. We run the experiments on the server with 16GB RAM, 50GB Disk SSD SATA 3, and 4 CPU Intel Core i5 (i5-8250U).
Table 14: Statistical information about prediction times (milliseconds) and size of classification models (Gigabyte) on the streaming system.
@@ 7 Conclusion and Future Work
In this paper, we constructed a new dataset called ViTHSD to detect targeted hate speech in Vietnamese texts. The dataset comprises 10,000 comments manually annotated by humans across five targets: Individual (1), Group (2), Religion/creed (3), Race/Ethnicity (4), and Politics (5). Each target includes four levels of hate: NORMAL, CLEAN, OFFENSIVE, and HATE, except NORMAL which indicates no specific target. We implemented a comprehensive annotation process, from annotator selection to consensus evaluation methodology. The ViTHSD dataset has a moderate level of inter-annotator agreement, measured by the Cohen-Kappa K-average metric ($k$=0.45). We assess multiple aspects of hate speech detection by testing several state-of-the-art models on this dataset. The XLM-R Bi-GRU-LSTM-CNN model achieves the best performance, with a Macro F1 score of 72.01% for the Target task only and 51.74% for the Target task with a level. Besides, the ViSoBERT - a monolingual pre-trained model on Vietnamese social texts, obtained better performance than XLM-R with fewer parameters and less time execution. Additionally, we have successfully constructed a streaming architecture that runs smoothly and demonstrated its functionality through a continuous 60-minute video stream, indicating the potential application of our work in the practical intelligence social listening systems to reduce hatred content.
Challenges such as vocabulary inconsistencies in social media texts, which include diverse languages, acronyms, slang, and mixed codes, result in inaccurate predictions. Addressing low-accuracy models will be a significant challenge going forward. Our error analysis revealed that developing a system capable of recognizing offensive language on social networks in Vietnam is incredibly challenging due to the complexity of the social texts created by users on social networks. Hence, our next studies will focus on improving the ability to understand the social texts of the model by normalizing the acronyms, slang, and code-mixing. Furthermore, we leverage the robust ability of the Large-language models (LLMs) to improve the performance of targeted hate speech detection tasks in the Vietnamese language.
@@
The source code and data are available for research purposes at: https://github.com/bakansm/ViTHSD